OpenStudy (dunk789):

Let f(x)=x^2-100. The average rate of change from x = -10 to x = 10 equals _____. A. 200 B. 100 C. 0 D. -100

OpenStudy (data_lg2):

\(\sf \Large Avg.\ Rate=\frac{f(10)-f(-10)}{10-(-10)}\) First, you have to find the values of f(10) and f(-10), then evaluate the formula above. \(\sf f(10)=(10)^2-100= ?\) \(\sf f(-10)=(-10)^2-100= ?\)